# Keyword-Aware ASR Error Augmentation for Robust Dialogue State Tracking

### v1

@@ 5 Conclusion
First, the introduction of keyword augmentation, while reducing "Wrong" and "Ignore" errors, resulted in an increase in "Spurious" errors, indicating a potential trade-off in error types. Additionally, our experiments were conducted using only the T5-base model. However, we believe that our method will work effectively with other models as well. Furthermore, we did not explore the application of our method to other similar datasets, such as those used in spoken language understanding (SLU) or named entity recognition (NER) tasks. Future work should consider these areas to validate and extend the applicability of our approach.
Appendix A Baseline Performance Comparison with Clean Text
Appendix B Training Details
@@
Appendix C Experiment Details
C.1 Dataset
  • Paraphrased environment: When recording the audio, the text was paraphrased to more closely resemble spoken language Soltau et al. (2022).
C.2 Comparison Methods
  • Back Translation Sennrich et al. (2015): We translated original texts to error texts and then back to the original texts for generating syntactic variations during the process. We use English to German 555facebook/wmt19-en-de and German to English666facebook/wmt19-de-en models as translator.
  • ASR translation: We employed a sequence-to-sequence structure to translate clean text into ASR-errored text. Our training set comprised 300 hours of paired clean and ASR-errored text, as detailed in Section 2.2. We fine-tuned the model based on the T5-base architecture Roberts et al. (2019), using the loss function defined in equation 4. The loss function is as follows:
    $L=-\sum_{i=1}^{I}\log P(e_{i}|g_{i}).$ (4)
Appendix D Prompt Example
Step 1 Prompt Example
Generate ASR error augmented text with similar pronunciation but different words based on the given gold text examples.
Error rate should be high
{Example 1}
Original: they have a single naupliar eye
ASR-errored: they have a single nor pure eye
{Example 2}
Original: i must have saint louis then huzza
ASR-errored: i must have st louis then hazard
{Example 3}
Original: i wonder uncle did not have her come
ASR-errored: i wonder uncle did not have a problem
{Example 4}
Original: i don’t want to have to pay for parking.
ASR-errored:
Step 2 Prompt Example
Change the key words in <hl> tag, to having a ASR error. ASR error has similar pronounciation with the correct word, but different charater.
Here is some example
{Example 1}
Original: I want to buy a book about <hl>luwombo best</hl> restaurant. Keywords : luwombo best
Result: I want to buy a book about luwambo vest restaurant.
{Example 2}
Original: hi, i’m looking for a bus that is depart from <hl>eliot<hl/> and arriving to <hl>holiday inn williamsport<hl/>? Keywords : eliot, holiday inn williamsport
Result: hi, i’m looking for a bus that is depart from Ellyot and arriving to holliday inn william’s port
{Example 3}
Original: the <hl>chabuton ramen<hl/> is a restaurant on the east. Keywords : chabuton ramen
Result: the shabuton raymond is a restaurant on the east.
{Example 4}
Original: i was hoping to see local faces while in cambridge. some <hl>entertainment</hl> would grate. Keywords: entertainment
Result:

### v2

@@ 4 Conclusion
Through detailed error analysis, we identified a trade-off introduced by our keyword-focused phonetic augmentation strategy. While the augmentation helps the model become more robust to noisy slot expressions—leading to substantial reductions in "Wrong" errors—it also increases the model’s sensitivity to phonetic variations. As a result, we observed cases where the model hallucinates slot values that were not actually mentioned, thereby increasing the number of "Spurious" errors. This hallucination effect represents a key limitation of our method. We attribute it to the model’s repeated exposure to noisy keywords, which may cause it to overgeneralize phonetic cues as valid slot mentions. As a direction for future work, we plan to incorporate an auxiliary loss term for slot presence prediction Heck et al. (2020); Kim et al. (2019) to help the model better distinguish between mentioned and unmentioned slots and mitigate this side effect.
Ethical Considerations
Our phonetic augmentation method, while effective for simulating ASR-style errors, may raise several ethical concerns. One such concern is the potential for accent bias, wherein phonetic transformations may disproportionately reflect majority or standard pronunciations, thereby marginalizing regional or minority accents. Another concern is the inadvertent corruption of proper names, particularly those that are less common or culturally specific, which could lead to misrepresentation or reduced inclusivity. We acknowledge these risks and emphasize that our method relies on LLMs trained on diverse and large-scale corpora. As such, the phonetic errors generated are likely to reflect dominant patterns present in mainstream ASR systems, rather than rare or region-specific variations. Nonetheless, we recognize the importance of fairness and inclusivity in language technologies and believe that future work should explore augmentation strategies that are more sensitive to accent and cultural variability.
Acknowledgements
This research was supported by the Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency funded by the Ministry of Culture, Sports and Tourism in 2025 (Project No. RS-2025-02413038, 45%), the IITP-ITRC (Information Technology Research Center) Program funded by the Ministry of Science and ICT (Project No. IITP-2025-RS-2024-00437866, 45%), and the IITP Artificial Intelligence Graduate School Program (POSTECH) funded by the Ministry of Science and ICT (Project No. RS-2019-II191906, 10%).
Appendix A Details of the EPA Method
A.1 Prompt Used for EPA
The prompts used in Step 1 and Step 2 are provided below.
Step 1 Prompt Generate ASR error augmented text with similar pronunciation but different words based on the given gold text examples.
Error rate should be high
. Example 1
Original: they have a single naupliar eye
ASR-errored: they have a single nor pure eye Example 2
Original: i must have saint louis then huzza
ASR-errored: i must have st louis then hazard Example 3
Original: i wonder uncle did not have her come
ASR-errored: i wonder uncle did not have a problem
. Now, following the above examples, generate an ASR-errored version of the following sentence:
Original: [Target utterance]
ASR-errored:
Step 2 Prompt Change the key words in <hl> tag, to having a ASR error. ASR error has similar pronounciation with the correct word, but different charater.
Here is some example.
. Example 1
Original: I want to buy a book about <hl>luwombo best</hl> restaurant.
Keywords : luwombo best Result: I want to buy a book about luwambo vest restaurant. Example 2
Original: hi, i’m looking for a bus that is depart from <hl>eliot<hl/> and arriving to <hl>holiday inn williamsport<hl/>?
Keywords : eliot, holiday inn williamsport Result: hi, i’m looking for a bus that is depart from Ellyot and arriving to holliday inn william’s port Example 3
Original: the <hl>chabuton ramen<hl/> is a restaurant on the east.
Keywords : chabuton ramen Result: the shabuton raymond is a restaurant on the east.
. Now, following the above examples, generate an ASR-errored version of the following sentence:
Original: [Target utterance with <hl> tag]
ASR-errored:
A.2 Detailed Keyword Highlighting Strategy for EPA
Example of adding <hl> tag
Original Hi, I need to go to Green Day hotel, then book a table at the Grill House.
Dialogue State hotel-name: Green Day, restaurant-name:Grill House
With <hl> tags Hi, I need to go to <hl>Green Day</hl> hotel, then book a table at the <hl>Grill House</hl>.
Table 7: Example of keyword highlighting using <hl> tags based on dialogue state annotations.
To explicitly introduce keyword-specific ASR errors, we first identify dialogue state values from the training corpus and match them against the user utterance ($\dot{u}$). Matched values are then automatically wrapped with <hl> tags based on slot annotations (e.g., DST slot labels or NER tags), as shown in Table 7. These highlighted utterances are passed to the LLM, which is instructed to perturb the text within the <hl> tags while preserving the rest. This keyword highlighting strategy is task-agnostic and can be easily applied to other keyword-sensitive tasks such as Named Entity Recognition (NER) or Spoken Language Understanding (SLU), where certain slot values or entities are critical for downstream prediction.
Example 1
Target can you tell me the address to the police station in point pleasant?
Retrieved 1 frayser station was not the depot on the point
 + ASR freya station was not the deep watch on the point
Retrieved 2 can you get me the maldeamores saga
 + ASR can you get me the melamorphos
Retrieved 3 cannot you tell her whom i am eh joseph
 + ASR cannot you tell her whom i am
Example 2
Target no, i just need to make sure it’s cheap. oh, and i need parking.
Retrieved 1 i need fifty ten foot long segments of wire
 + ASR i need fifty ten foot long signals of my life
Retrieved 2 a drive with a different encoding mechanism would need different patterns
 + ASR and drive was a different building recognition would need different patterns
Retrieved 3 to reach to calcutta you need less time to reach dhaka
 + ASR to reach tukaukara you need last time to reach daka
Example 3
Target i’m open to any kind of food. i’m looking for something in the centre and on the expensive side.
Retrieved 1 kokai means open to the public or laid open
 + ASR cook eye means open to the public all laid open
Retrieved 2 the town of beauharnois was the major centre
 + ASR the town of bo hanwa was the major center
Retrieved 3 the gate is open at eleven
 + ASR the gate is open at 11
Table 8: In-context examples retrieved based on phoneme-level similarity. For each target utterance (top row), we retrieve three ($g$, $e$) example pairs from the database using phonetic similarity between the target and $g$. Colored segments highlight phonetically similar phrases between the target and retrieved examples.
A.3 Retrieved In-Context Example
In Section 2.2, we retrieved in-context examples based on phoneme-level similarity. Table 8, we present several representative examples to illustrate this retrieval process, showing how phonetically similar phrases (highlighted in color) are matched between the target and retrieved utterances. This demonstrates that the retrieval mechanism effectively captures pronunciation-level patterns relevant to ASR-style errors.
A.4 Additional Examples of ASR-style Errors
Original Augmented
- no, i just need to make sure it’s cheap. oh, and i need parking - no, i just need to make sure it’s sheep. oh, and i need parking.
- i am departing from marion - - i am departing from maryland
- no, i don’t need anything else right now. thank you for your assistance. good bye. - no, i don’t need anyone else right now. thank you for your persistence. good buy.
- i would like to go to sandy please. - i would like to go to cindy please.
- i would like to keep it in the moderate range, please. - i would like to keep it in the mod rain, please
- could i get the address for it? i would also like an expensive place to eat around it. - could i get the actress for it? i would also like an extensive place to eat around it.
- i need to take a train out of garrett, i will be leaving town on wednesday - i need to make a plane out of garrett, i will be weaving town on wednesday.
- do you have any indian restaurants in the south in a different price range? - do you have any indonesian restaurants in the south in a different prize range
- nope, same people. - nope, same pupil.
- i’m looking for a college type attraction - i’m looking for a knowledge-type action.
- yes, please book me a room for friday - yes, please cook me a broom for friday
- yes, could you please email or fax me the fare amount, as well as the reference number? - yes, could you please email or text me the fair amount, as well as thereference code?
- ois el shaddai a guest house or hotel? iz let shadai a gest house or motel?
- great! i also need a train from mount pleasant to sabattus, please. - great! i also need a strain from mount pleasant to suspicious, please
- yes, can you help me find a train that can take me from lovelock to abbot? - yes, can you help me find a plane that can take me from love lock to rabbit?
Table 9: Examples of augmented utterances generated by injecting phoneme-level ASR-style errors. For each original utterance (left), the corresponding augmented version (right) includes substitutions that mimic realistic ASR recognition mistakes. Blue-colored phrases indicate changes in keywords that are used as slot values in DST, while orange-colored phrases represent overall ASR-style errors.
Table 9 presents additional examples of ASR-style errors generated by our EPA method, including both general and keyword-specific transformations.
Appendix B Experimental Setup
B.1 Details of the ASR Environment
  • Paraphrased environment: When recording the audio, the text was paraphrased to resemble more natural, real-life spoken languageSoltau et al. (2022).
B.2 Comparison Methods
  • Back Translation Sennrich et al. (2015): We translated original texts to error texts and then back to the original texts for generating syntactic variations during the process. We use English to German 333facebook/wmt19-en-de and German to English444facebook/wmt19-de-en models as translator.
  • ASR translation: We employed a sequence-to-sequence structure to translate clean text into ASR-errored text. Our training set comprised 300 hours of paired clean and ASR-errored text. We fine-tuned the model based on the T5-base architecture Roberts et al. (2019), using the loss function defined in equation 3. The loss function is as follows:
    $L=-\sum_{i=1}^{I}\log P(e_{i}|g_{i}).$ (3)
B.3 Training Details
@@
Appendix C Further Experiments
C.1 Statistical Significance Analysis
Method Low-acc ASR Noised Aud. Paraphrased High-acc ASR
JGA N-acc JGA N-acc JGA N-acc JGA N-acc
Baseline
AEDA ns ns ns ns ns ns ns ns
EDA * * ** ** ns ** ** **
BT ** ** ** ** * ** ** *
TTS-ASR ns ns ns ns ns ns ns ns
ASR trans. ns * ns ns ns * * *
EPA (GPT-3.5) *** ** *** *** ** ** * **
Table 10: Statistical significance results compared to the Baseline using paired t-tests across three random seeds. Stars indicate significance levels: * for $p<0.05$, ** for $p<0.01$, *** for $p<0.001$, and ns for non-significant differences.
To assess the reliability of our results, we conducted paired t-tests between each method and the Baseline to determine whether the observed performance improvements are statistically significant. We report 95% confidence intervals to reflect performance variability. Statistical significance is denoted using asterisks: $\ast$ for $p<0.05$, $\ast\ast$ for $p<0.01$, and $\ast\ast\ast$ for $p<0.001$.
As shown in Table 10, EPA (GPT-3.5) achieves statistically significant gains in nearly all evaluation settings, particularly under low-accuracy and noised ASR conditions. These results confirm that the improvements brought by our method are both consistent and statistically reliable.
C.2 Baseline Performance Comparison with Clean Text
C.3 Effect of Keyword Augmentation
Method Weak-ASR Noised Audio Strong ASR
JGA N-acc JGA N-acc JGA N-acc
Baseline 29.88 45.76 29.70 46.77 34.87 52.07
+ Overall Error (Step 1) 31.31 50.67 31.13 52.29 35.40 55.80
+ Keyword Error (Step 2) 31.66 49.88 31.61 51.02 36.12 54.23
EPA (Step1+2) 32.39 51.12 32.24 52.70 36.61 55.87
Table 11: Ablation on overall and keyword-specific augmentation steps. Combining both (EPA) yields the highest robustness across ASR conditions.
To further validate the contribution of each component in our method, we conducted an ablation study separating Step 1 (overall error augmentation) and Step 2 (keyword-specific error augmentation), using GPT-3.5 as the generator. Both augmentation strategies individually improved performance over the baseline. Step 1 yielded consistent gains across all ASR settings, while Step 2 notably enhanced slot-value accuracy. When combined (EPA, ours), the model achieved the highest robustness in all evaluation conditions, confirming that the two strategies are complementary.
C.4 Experiments with a Different Baseline
Method Low-acc ASR Noised Aud. Paraphrased
JGA N-acc JGA N-acc JGA N-acc
Baseline 29.9 45.82 27.25 41.81 25.81 44.51
   + EPA 30.63 48.54 27.5 46.52 27.65 47.96
Table 12: Experiment with GPT-2 model as baseline.
In the main experiments, we use T5-base as the backbone model. To assess the generalizability of our approach, we additionally conduct experiments using a GPT-2 Radford et al. (2019) model, as shown in Table 12. The results show a consistent trend with those of T5-base, demonstrating that our method is effective across different backbone architectures.
C.5 Generalizability to Other Tasks
Method Low-acc ASR Noised Audio High-acc ASR
Baseline 56.29 60.50 60.02
+ OPT 6.7B 59.64 62.84 62.47
+ LLaMA 7B 58.29 60.53 60.46
+ GPT-3.5 (125B) 59.39 62.62 62.16
Table 13: NER results on the ASAPP/SLUE dataset under different ASR conditions. We reported the F1 score.
Method Low-acc ASR Noised Audio High-acc ASR
Baseline 55.25 63.23 64.32
+ OPT 6.7B 56.97 64.94 66.10
+ LLaMA 7B 57.24 64.26 65.56
+ GPT-3.5 (125B) 58.78 66.16 67.49
Table 14: SLU results on the SLURP dataset under different ASR conditions. We reported F1 score.
To evaluate the generalizability of our approach beyond the DST domain, we extended our experiments to two additional spoken language understanding tasks: Named Entity Recognition (NER) and Spoken Language Understanding (SLU). We applied our EPA methodology under three ASR conditions—low-accuracy ASR, noised audio, and high-accuracy ASR—using the same experimental setup as in the DSTC11 evaluation. We used asapp/slue dataset for NER taskShon et al. (2022), and SLURP dataset Bastianelli et al. (2020) for SLU task.
The results, shown in Table 13 and Table 14, demonstrate that our method consistently improves performance across all ASR conditions for both NER and SLU tasks. Notably, the gains are especially prominent under low-accuracy and noisy conditions, confirming that our approach is broadly applicable to other tasks.
C.6 Additional Fine-grained Metrics
Method Precision Recall F1 Slot Accuracy
Baseline 50.5 50.1 50.3 92.8
+ TTS-ASR 49.5 49.6 49.5 93.4
+ ASR-Translation 51.8 51.7 51.7 93.1
+ EPA (GPT3.5) 55.2 54.8 55.0 93.7
Table 15: Fine-grained DST metrics under Low-accuracy ASR setting.
Method Precision Recall F1 Slot Accuracy
Baseline 52.0 51.3 51.6 92.5
+ TTS-ASR 53.7 53.1 53.4 92.8
+ ASR-Translation 53.8 53.0 53.4 92.5
+ EPA (GPT3.5) 56.1 55.3 55.7 93.2
Table 16: Fine-grained DST metrics under Noised ASR setting.
To supplement the main results focusing on JGA and N-Acc, we report additional fine-grained metrics—Precision, Recall, F1, and Slot Accuracy—under two ASR corruption settings: Low-accuracy ASR and Noised ASR. These metrics provide a more comprehensive view of model behavior in diverse error conditions in table 15 and  16.
C.7 Results with Different Random Seeds
@@
Table 17 reports the results of our main experiments (Table 2) repeated with three different random seeds, demonstrating the consistency of the observed trends.
Appendix D Details of Quality Evaluation
D.1 About Metric
As described in Section 3.4, we use a phonetic similarity metric to evaluate pronunciation-level consistency between the original and augmented text. Specifically, we compute the normalized phoneme edit distance, which quantifies the minimal number of phoneme-level operations required to transform one utterance into another. A higher score indicates greater phonetic similarity. We used the eng-to-ipa library [^5] https://pypi.org/project/eng-to-ipa/ for phoneme conversion in our implementation, as shown in the code snippet below.
def phonetic_similarity(original_text, augmented_text):
original_ipa = to_phoneme(original_text)
augmented_ipa = to_phoneme(augmented_text)
edit_distance = nltk.edit_distance(original_ipa, augmented_ipa)
# Normalize the edit distance
max_length = max(len(original_ipa), len(augmented_ipa))
normalized_distance = float(edit_distance) / float(max_length)
# Convert to similarity score
similarity_score = 1 - normalized_distance
return similarity_score
D.2 Human Evaluation Details
To assess the plausibility of the generated ASR-style errors, we conducted a human evaluation involving two graduate students. Each participant was asked to rate whether a given sentence transformation could plausibly be attributed to an ASR error, using a 4-point Likert scale:
  • 1 – Not at all: The change is unlikely to be due to an ASR error. It appears to stem from other factors such as meaning alteration or stylistic variation.
  • 2 – Unlikely: The transformation is probably not caused by an ASR error.
  • 3 – Somewhat likely: The transformation may plausibly be caused by an ASR error.
  • 4 – Very likely: The transformation clearly appears to result from an ASR error.
Each sentence pair (original and transformed) was rated independently by both annotators. Inter-rater agreement and average scores are reported in Section 3.4. The distribution of Likert scores for each annotator is as follows : Annotator 1 assigned 5% of scores as 1, 7% as 2, 17% as 3, and 71% as 4. Annotator 2 assigned 3% of scores as 1, 24% as 2, 53% as 3, and 20% as 4.